OpenStudy (anonymous):

4). A baseball player gets 4 hits in the first 12 games of the season. If he continues hitting at the same rate, how many hits will he get in the first 48 games?

OpenStudy (anonymous):

16. 48/12=x/4 . 12x=192 x=16